An electrical firm which manufactures a certain type of bulb wants to estimate its mean life. The life of the light bulb is normally distributed with a standard deviation of 40 hours. A random sample of 36 bulbs resulted in a mean of 200 hours.

a) (3 points) Construct a 92% confidence interval for the mean life of all light bulbs the firm manufactures.

b) (4 points) How many bulbs should be tested so that we can be 92% confident that the estimate of the mean will not differ from the true mean life by more than 10 hours?
Answer #1


a) The sample mean is \overline{x}=200 , the sample standard deviation is \sigma =40 and sample size is n=36 .

The (1 - \alpha ) 100\% confidence interval for true mean based on the sample mean is

\overline{x} \pm z_{1-\alpha /2}\frac{\sigma }{\sqrt{n}}\\

The 92% CI for mean \left (\alpha =0.08 \right ) is

200 \pm z_{1-0.08 /2}\frac{40}{\sqrt{36}}\\ {\color{Blue} (188.33, 211.67)}

a) We need

z_{1-0.08 /2}\frac{40}{\sqrt{n}}<10\\ n> z^2_{1-0.08 /2}40^2/10^2=49.03843

The required sample size is {\color{Blue} 50} .